OpenStudy (anonymous):

Find an equation for the nth term of the arithmetic sequence. -1,2,5,8 ...

OpenStudy (anonymous):

Add three to each sequence

OpenStudy (anonymous):

here are my answer choices an=-1+3(n) an=-1+3(n-1) an=-1+3 an=-1+3(n+1)

OpenStudy (cggurumanjunath):